OpenStudy (anonymous):

give a formula for the solution y(x) of the differential equation y'(x) = 0.4y(x) with y(0)=8

OpenStudy (anonymous):

hmm...a separable differential equation\[ y'(x) = 0.4y(x)\]\[\frac{\text{d}y}{\text{d}x}=0.4y\]\[\frac{\text{d}y}{y}=0.4\ \text{d}x\]

OpenStudy (anonymous):

but what is the y(x) function?

OpenStudy (anonymous):

i got y(x) =e^(0.4(0.5x^2+20)) but its incorrect

hartnn (hartnn):

first,did u integrate that? dy/y=0.4x what did u get? write that

OpenStudy (anonymous):

i got y'(x)/0.4y(x) = 1

OpenStudy (anonymous):

this is all steps i wrote it, but it still incorrect

hartnn (hartnn):

this is simpler than u solved,u just need to integrate this on both sides, \[\frac{dy}{y}=0.4dx\]

OpenStudy (anonymous):

someone suggest me write as this ln(y) = 0.4x +c then y(0)=8 will gives me ln(8)0=c. therefore, ln(y)=0.4x+ln(8))

OpenStudy (anonymous):

is that the same thing as yours?

hartnn (hartnn):

after integrating that u WILL get that: as \[\int\limits_{}^{}dy/y=\ln y\] and \[\int\limits_{}^{}0.4dx=0.4x+c\] so yes,same thing

OpenStudy (anonymous):

i got the lny(x) = 2x/5 + 8

hartnn (hartnn):

but how did u get c as 8?

OpenStudy (anonymous):

c is ln(8)

OpenStudy (anonymous):

y= e^(2x/5 + ln(8))

hartnn (hartnn):

now thats correct :)

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

also can you help me on this one?

hartnn (hartnn):

sure.

OpenStudy (anonymous):

i did several times, still incorrect. i got the tips that need use integral by part, u = sinIntegral(x)

hartnn (hartnn):

this seems difficult for me too,need some time.

OpenStudy (anonymous):

i think we don't need take the sinIntegral(x) apart. we could see it as a whole thing?

hartnn (hartnn):

i have never solved such integral.....this reference gives the integral of sinintegral x: http://mathworld.wolfram.com/SineIntegral.html

OpenStudy (anonymous):

emm ok. thx for taking ur time

hartnn (hartnn):

i would hand this back to @mukushla